OpenStudy (zyberg):

Question about binomial theorem is in comments (due to a need of correct markup).

OpenStudy (holsteremission):

This isn't necessarily true. Consider \(n=1\) and \(k=0\).

OpenStudy (zyberg):

\[\left(\begin{matrix}n \\ k\end{matrix}\right)<\left(\begin{matrix}n \\ k+1\end{matrix}\right)\] if and only if 0 <= k < 1/2(n-1)

OpenStudy (zyberg):

Sorry, @HolsterEmission, it seems that another post of mine, where I tell that it is 0 <= k < 1/2(n-1) did not send.

OpenStudy (welshfella):

n! n! ---- < ------- k!(n-k)! (k+1)! (n - k -1)! k!(n-k)! > k+1)! (n - k -1)! we should be able to get a relation between k and n from this.

OpenStudy (zyberg):

@welshfella should we take the base as k or as n? This time I am confused because of two letters...

ganeshie8 (ganeshie8):

Hey, what's the question ?

OpenStudy (welshfella):

I'm taking it to be Zybergs first post...

OpenStudy (welshfella):

- proof of that inequality

ganeshie8 (ganeshie8):

Oh I see... nice :)

ganeshie8 (ganeshie8):

k!(n-k)! > (k+1)! (n - k -1)! simplifying above should do yeah

OpenStudy (welshfella):

k!(n-k)! > (k+1)! (n - k -1)! k!(n-k)! > k!(k+1) (n-k-1)! (n - k)! > (k+1)(n - k - 1)! Now right the left side as (n-k)(n - k - 1)! and you are nearly done.

OpenStudy (zyberg):

Oh, so it won't even require real induction? The thing that is confusing me is that I usually have to use induction on theorems that have only one unknown...

OpenStudy (zyberg):

@ganeshie8 it's from "Elementary Numbers Theory" book, you recommended to me ;)

ganeshie8 (ganeshie8):

I think we may use induction by inducting on "n"

ganeshie8 (ganeshie8):

You may start by proving the given statement is true for n=3

ganeshie8 (ganeshie8):

Notice that we're not interested in n < 3 because there is not integer "k" that satisfies the given condition 0 <= k < 1/2(n-1) when n < 3.

OpenStudy (zyberg):

Yes, I tried to do that, but then I need to set up a value for k too, yes? How would that work with induction step then? (I understand setting base to be n = 3, k = 0)

ganeshie8 (ganeshie8):

You don't have to worry about "k" at all as we're inducting on "n"

ganeshie8 (ganeshie8):

Let me say again, we're inducting on "n", not k.

OpenStudy (zyberg):

But placing the value of n = 3 would get us: (3-k)! > (k+1)(2-k)!, how would we check, if it's true then?

OpenStudy (zyberg):

Without knowing k, is it possible to rework (2-k)! somehow, to prove to be true?

ganeshie8 (ganeshie8):

use the given condition : 0 <= k < 1/2(n-1) plugin n=3 and see what values of "k" we must check

OpenStudy (zyberg):

But wouldn't that account as "worrying" for k? k would become 0, but what about induction step? What k would be for n = r? n = r+1?

ganeshie8 (ganeshie8):

We are still not worrying about "k" as part of the induction process : step1) show the statement is true for n=3 step2) assume that the statement is true for n = r step3) prove the statement is trur for n = r+1 whenever step2 is true.

ganeshie8 (ganeshie8):

For n=r, we assume that the given statement is true for all "k" that satisfy 0 <= k < 1/2(r-1)

ganeshie8 (ganeshie8):

We just have to follow the rules when using induction to prove

ganeshie8 (ganeshie8):

Proof happens pretty much automatically

OpenStudy (zyberg):

so: (t - k)! > (k+1)(t - k - 1)! Multiplying by t - k + 1 or something like that?

ganeshie8 (ganeshie8):

Yep, also note that we're only proving below part : If 0 <= k < 1/2(n-1), then \(\left(\begin{matrix}n \\ k\end{matrix}\right)<\left(\begin{matrix}n \\ k+1\end{matrix}\right)\).

ganeshie8 (ganeshie8):

We need to prove the converse of that too

OpenStudy (welshfella):

Ah yes that is what I was attempting to do - the converse.

OpenStudy (welshfella):

i misread the question really

OpenStudy (zyberg):

You are right about that... wouldn't have thought of it ;) But how would we prove converse of that?

OpenStudy (welshfella):

I think if you further simplify: (n - k)! > (k+1)(n - k - 1)! (n-k)(n-k-1)! > (k+1)(n ik-1)! n - k > k + 1 k < (1/2)(n- 1)

OpenStudy (welshfella):

and therefore 0 <= k < 1/2(n-1)

OpenStudy (zyberg):

Okay, thank you very much! :) I think that I need a break now, for not noticing that... Ehh! Thank you for help as well, @ganeshie8! Decided to finish the book until the end of the year and so far it's quite fun (though I have been jumping all around of it and only now I started doing things systematically).

OpenStudy (welshfella):

yw I'm going for a break too!

ganeshie8 (ganeshie8):

Its my favorite book on NT. Good to see you're enjoying it too :)